OpenStudy (anonymous):

solve the linear equation x=13-2y y=x-8

OpenStudy (shamil98):

substitution. let's solve for y. y = 13 -2y -8

OpenStudy (anonymous):

\[x = 13 - 2(x-8)\] You can find x, then find y. :D

OpenStudy (shamil98):

3y =5 y = 5/3 input that into the equation for x to find your x value.

OpenStudy (anonymous):

@SACAPUNTAS i did it like this y=13-2(x-8) y=13-2x+16 y=29+ -2x y= -14.5 what now? cx

OpenStudy (anonymous):

You wrote y but you meant x, right? Otherwise that doesn't make a lot of sense. I am with you up to: \[x = 29 - 2x\]But then since it's x you end up with: \[x = \frac{29}{3}\]

OpenStudy (anonymous):

@SACAPUNTAS yeah, my bad why do you end up with 3 and not 2?

OpenStudy (anonymous):

Because you have to add 2x to both sides to get rid of the -2x. \[x = 29 - 2x\]\[x + 2x = 29 - 2x + 2x\]\[3x = 29\]

OpenStudy (anonymous):

is there an easier way to do that? like i thought you just had to divide it

OpenStudy (anonymous):

There might be an easier way but I don't know it. :D

OpenStudy (anonymous):

oh okay cx lol thank you so much!